OpenStudy (anonymous):

cos^2 x + sinx+1 = 0

myininaya (myininaya):

\[\text{ use } \cos^2(x)=1-\sin^2(x)\]

myininaya (myininaya):

Put it in terms of one trig function By using the identity I mentioned above you can do this

OpenStudy (anonymous):

o okay so all cosine

myininaya (myininaya):

\[\text{ since } \cos^2(x)=1-\sin^2(x) \text{ then you can replace } \cos^2(x) \text{ with }\] \[1-\sin^2(x)\]

myininaya (myininaya):

So it will be in terms of sine not cosine

OpenStudy (anonymous):

oh ok can u help me with something else while ur here

myininaya (myininaya):

\[1-\sin^2(x)+\sin(x)+1=0\] Do you see that I just replaced \[\cos^2(x) \text{ with } 1-\sin^2(x)\]

OpenStudy (anonymous):

yaaa i remmeber tha rule from class

OpenStudy (anonymous):

ok so if i have cos x = -0.438

OpenStudy (anonymous):



myininaya (myininaya):

So this is a different question?

OpenStudy (anonymous):

is that the reference angle?

OpenStudy (anonymous):

ya

myininaya (myininaya):

So you want to find x right?

OpenStudy (anonymous):

wait so do i type it in calc and that ALWAYS gives me the angle in quadrant 1?